# MoA: Mixture of Sparse Attention for Automatic Large Language Model Compression

### v1

@@ B.1.2 Long-Context Retrieval
We conduct a detailed experiment to test the retrieval ability of different attention methods across various attention spans and input lengths. Following previous work [7, 60], we quantify effective context length as the maximum input length where retrieval accuracy remains above a 90% threshold. Figure 6 shows the detailed data for effective context length calculation. As shown in the figure, StreamingLLM can hardly maintain retrieval accuracy when the input length is beyond the attention span, while MoA can effectively extend the effective context length.
@@ B.1.3 Long-Context Understanding
We conduct experiments with various densities on the LV-Eval benchmark [70]. As shown in Figure 7, MoA constantly outperforms the uniform static attention baseline StreamingLLM at various densities, demonstrating the effectiveness of our heterogeneous elastic rules.
@@ B.2.1 Throughput Breakdown
Table 7: Improvements in throughput (tokens per second) across various configurations and input sizes for MoA over FlashAttn2 on models with 7b and 13b parameters. The batch sizes for FlashAttn2 are configured to the maximum capacity of an A100-80GB GPU.
4k 8k 16k
Size Configuration Batch Throughput Batch Throughput Batch Throughput
7B FlashAttn2 30 134.6 15 66.9 8 32.9
+Static KV-Cache 30 496.1 15 219.5 8 91.6
+Reduced Attention 30 722.5 15 369.9 8 178.3
+Increased Batch Size 50 897.7 25 436.7 12 206.4
13B FlashAttn2 16 81.3 8 40.8 4 19.8
+Static KV-Cache 16 264.6 8 111.3 4 62.2
+Reduced Attention 16 329.6 8 156.4 4 87.3
+Increased Batch Size 28 471.5 14 222.6 7 108.3
In this section, we conduct a detailed analysis of the throughput breakdown for MoA, compared to the baseline comprising Huggingface with FlashAttention. The observed increase in throughput, as shown in Table 7, primarily stems from three aspects:
@@
B.2.2 Comparison with vLLM
Table 8: Throughput comparison of MoA and vLLM on 7B and 13B models.
Throughput (tok/s)
Size Framework 4k 8k 16k
7B vLLM 628.8 323.0 145.5
MoA 897.7 436.7 206.4
13B vLLM 314.8 160.5 71.1
MoA 471.5 222.6 108.3
We compare the throughput of MoA with the highly optimized inference framework vLLM [33] as shown in Table 8. MoA still achieves a 1.4×1.4$\times$1.4 × to 1.5×1.5$\times$1.5 × increase in throughput. Notably, vLLM incorporates extensive system-level optimizations that are orthogonal to MoA, suggesting the potential for MoA to further improve. As vLLM constantly pre-allocates a 71GB memory and automatically manages the actual batch size, we exclude memory metrics from the comparison.
B.2.3 Automatic Optimization Overhead
Table 9: Time overhead for various stages of MoA.
Stage Time
Calibration Dataset Generation 15 min
Profile 25 min
Optimize 25 min
Validate 40 min
Sum. 105 min
@@ B.3.1 Calibration Dataset
Table 10: Performance comparison on various test sets, using different calibration sets. Tested on Vicuna-7B model. The result is tested with 50% density MoA on LongBench [4] 0-4k split.
@@
In this section, we validate the robustness of our calibration dataset design principles. We select three sub-tasks and respective datasets from the LongBench benchmark, including Qasper [15], MultiNews [17], and TREC [38, 28]. We use their training set to construct the calibration dataset, and use their respective test set in LongBench to calculate the score. Following Section 5, all calibration datasets are constructed using the original model’s response to the context and questions as the supervision.
@@ C.1 Statistics on Rules Discovered by MoA
This subsection provides empirical evidence for rules discovered by MoA as mentioned in Section 6.4. The lines and spans in Figure 8 show that all heads at the first few layers generally need a high KV-Cache density. Following that, a few layers generally only require medium density. Then, in the final layers, most heads require low density, while some outlier heads need high density. This observation conforms to previous findings of the intrinsic dimension of LLM [61]. The geometry of density is similar to the intrinsic dimension of LLM, with two local minima. As observed in Figure 8, layers with lower average density (smaller values on the lines) typically display a wider range of density (wider shades). Figure 9 validates such observation. This observation confirms the need for heterogeneous attention rules within the same layer.
@@ C.2.1 Mask Visualization and Semantic Categorization
Given any token, two kinds of information are used as the model inputs: position encoding and token embedding. Position encoding indicates the absolute [74] or relative positions [59] of tokens in the sentence. Token embedding maps different tokens as different vectors. The attention head $h$ responds to both information and output the corresponding attention value $A_{h}$. As shown in equation 6, we denote the influence of position and token of head $h$ as function $P_{h}$ and $T_{h}$, respectively. The attention value $A_{h,i,j}$ between the $i$th and $j$th token $t_{i}$ and $t_{j}$ is determined by the combination $f_{h}$ of position and token influence functions.
@@ C.2.2 Quantitative Semantic Analysis
Appendix D MoA Automatic Pipeline Details
D.1 Derivation of Attention Influence
We use the first-order Taylor expansion to calculate the influence of each attention value. This approximation approach is supported by methodologies commonly employed in other LLM compression approaches [36, 55, 14, 29].
@@ D.1 Derivation of Attention Influence
D.2 Optimization Details
D.2.1 Optimizing at Single Length
@@
D.2.2 Optimizing at Multiple Lengths
With the ability to optimize at a single length, we utilize the same framework for multi-objective MIP across various lengths. The key is to transform the multi-objective MIP problem into several single-objective MIP problems [47]. We utilize the idea of epsilon-constraint method [72].

### v2

@@ B.1.2 Long-Context Retrieval
Figure 7: Retrieval accuracy tests on LongEval with Vicuna-7B. (a) Varies input lengths and densities to show effective context lengths across attention spans, (b) Set input length at 8k and show retrieval accuracy across different densities.
We conduct a detailed experiment to test the retrieval ability of different attention methods across various attention spans and input lengths.
Figure 6 shows the detailed data for effective context length calculation. As shown in the figure, StreamingLLM can hardly maintain retrieval accuracy when the input length is beyond the attention span, while MoA can effectively extend the effective context length.
Following previous work (Chen et al., 2023; Tworkowski et al., 2023), we quantify effective context length as the maximum input length where retrieval accuracy remains above a 90% threshold. As shown in Figure 7(a), StreamingLLM and H2O achieve effective context lengths of no more than 2k tokens beyond their attention spans. In contrast, MoA expands its effective context length to approximately 3.9×3.9$\times$3.9 × its attention span before reaching up to the 12k limit of the original model. Figure 7(b) further shows that at a fixed input length of 8k, MoA reaches over 0.9 retrieval accuracy with just 25% density, whereas StreamingLLM and H2O require 100% and 75% density, respectively.
@@ B.1.3 Long-Context Understanding
We conduct experiments with various densities on the LV-Eval benchmark (Yuan et al., 2024). As shown in Figure 8, MoA constantly outperforms the uniform static attention baseline StreamingLLM at various densities, demonstrating the effectiveness of our heterogeneous elastic rules.
Table 7: LongBench scores for various models with different attention methods. All methods employ 50% density in the decode stage.
LongBench $\uparrow$
Model Attention 0-4k 4-8k 8-16k
Vicuna-7B Original 37.91 33.82 32.54
H2O 36.23 32.74 31.81
InfLLM 35.23 33.54 30.15
StreamingLLM 30.53 33.28 31.70
MoA 37.04 32.90 31.94
Vicuna-13B Original 42.25 39.52 35.93
H2O 41.63 38.02 34.75
InfLLM 39.36 37.66 34.36
StreamingLLM 30.65 33.07 32.68
MoA 41.73 38.88 35.69
Llama3-8B Original 44.27 43.53 43.26
H2O 43.46 43.01 42.50
InfLLM 42.78 42.69 41.81
StreamingLLM 37.20 38.02 39.43
MoA 43.07 42.75 43.09
Llama3-70B Original 50.70 48.05 48.55
H2O 50.16 47.77 OOM
StreamingLLM 45.14 42.40 40.04
MoA 49.74 46.80 46.84
Table 8: Performance comparison across different models and attention methods with the LV-Eval dataset. The numbers in brackets indicate the number of sub-datasets for the category.
Single-QA Multi-QA Retrieval
Model Attention w/o. Conf (2) w. Conf (2) w/o. Conf (3) w. Conf (2) w. Conf (2)
Vicuna-7B Original 10.49 6.29 6.83 5.60 0.00
H20 9.16 6.20 6.44 4.80 0.00
InfLLM 7.11 6.70 6.07 4.80 0.00
StreamingLLM 7.54 5.90 5.98 3.56 0.00
MoA 9.98 6.27 6.16 5.31 0.09
Vicuna-13B Original 10.64 7.28 5.32 5.07 1.08
H20 9.53 6.54 5.25 5.36 1.83
InfLLM 10.21 9.35 6.03 3.19 2.08
StreamingLLM 9.05 5.86 5.37 3.19 3.70
MoA 11.04 6.93 5.79 5.84 6.88
Llama3-8B Original 34.05 19.51 11.41 17.70 7.84
H20 28.52 17.05 11.11 15.98 9.95
InfLLM 24.94 17.75 10.61 14.80 6.04
StreamingLLM 20.21 9.57 8.14 9.36 10.03
MoA 32.98 20.53 10.65 17.57 8.98
Llama3-70B Original 44.44 25.02 16.71 22.86 17.43
StreamingLLM 26.63 14.22 14.04 14.70 19.38
MoA 42.44 23.58 15.75 21.27 19.19
Figure 9: (a) LV-Eval and (b) LongBench scores for different attention methods at 50% density, tested on Vicuna-7B, 13B and Llama3-70B models. Scores normalized against the original dense model.
We detailed the respective scores for LongBench and LV-Eval in Table 7 and Table 8. The number in the bracket of Table 8 indicates the number of sub-datasets for the category.
@@ B.2.1 Throughput Breakdown
Table 9: Efficiency analysis of different frameworks on 7B and 13B models. H2O and MoA use 50% density. GPU memory evaluated with batch sizes 8 (7B model) and 4 (13B model).
Memory (GB)
Size Framework 4k 8k 16k
7B FlashAttn2 28.5 44.4 76.3
H2O 36.9 OOM OOM
MoA 22.7 32.9 53.5
13B FlashAttn2 36.8 49.2 74.0
H2O 40.4 77.9 OOM
MoA 32.0 39.6 55.0
In this section, we explain the decode throughput breakdown for MoA, compared to the baseline comprising Huggingface with FlashAttention2. The observed increase in throughput, as shown in Table 5, primarily stems from three aspects:
@@
B.2.2 Automatic Compression Pipeline Overhead
Table 10: Compression overhead for various stages of MoA across models with differing parameter sizes, reported as the amount of GPU $\times$ latency, except when only one GPU is used. Larger models necessitate more GPUs due to model parallelism. All stages utilize GPUs, except for the Optimize stage, which uses the CPU.
Stage 7B LLM 13B LLM 70B LLM
Calibration Data Gen. 10min 15min 2 $\times$ 60min
Profile 20min 2 $\times$ 25min 8 $\times$ 210min
Optimize (CPU) 30min 25min 100min
Validate 35min 40min 2 $\times$ 140min
Total Latency 1h 35min 1h 45min 8h 30min
Total GPU Time 1h 5min 1h 45min 34h 40min
Table 11: Progressive compression overhead for various stages of MoA with respect to different parameter sizes and calibration (validation) dataset sizes.
Stage Complexity w.r.t parameter size Complexity w.r.t dataset size
Calibration Dataset Gen. Linear Linear
Profile Linear Linear
Optimize Polynomial $\sim$ Exponential for #Head Irrelevant
Validate Linear Linear
Empirical Latency Almost Linear Linear
@@ B.2.2 Automatic Compression Pipeline Overhead
We also show the progressive compression overhead for MoA in Table 11.
@@ B.3.1 Calibration Dataset
Table 12: Performance comparison on various test sets, using different calibration sets. Tested on Vicuna-7B model. The result is tested with 50% density MoA on LongBench (Bai et al., 2023) 0-4k split.
@@
In this section, we validate the robustness of our calibration dataset design principles. We select three sub-tasks and respective datasets from the LongBench benchmark, including Qasper (Dasigi et al., 2021), MultiNews (Fabbri et al., 2019), and TREC (Li & Roth, 2002; Hovy et al., 2001). We use their training set to construct the calibration dataset, and use their respective test set in LongBench to calculate the score. Following Section 5, all calibration datasets are constructed using the original model’s response to the context and questions as the supervision.
@@ C.1 Statistics on Rules Discovered by MoA
This subsection provides empirical evidence for rules discovered by MoA as mentioned in Section 6.5. The lines and spans in Figure 10 show that all heads at the first few layers generally need a high KV-Cache density. Following that, a few layers generally only require medium density. Then, in the final layers, most heads require low density, while some outlier heads need high density. This observation conforms to previous findings of the intrinsic dimension of LLM Valeriani et al. (2023). The geometry of density is similar to the intrinsic dimension of LLM, with two local minima. As observed in Figure 10, layers with lower average density (smaller values on the lines) typically display a wider range of density (wider shades). Figure 11 validates such observation. This observation confirms the need for heterogeneous attention rules within the same layer.
@@ C.2.1 Mask Visualization and Semantic Categorization
Given any token, two kinds of information are used as the model inputs: position encoding and token embedding. Position encoding indicates the absolute (Zhang et al., 2022) or relative positions (Touvron et al., 2023) of tokens in the sentence. Token embedding maps different tokens as different vectors. The attention head $h$ responds to both information and output the corresponding attention value $A_{h}$. As shown in equation 6, we denote the influence of position and token of head $h$ as function $P_{h}$ and $T_{h}$, respectively. The attention value $A_{h,i,j}$ between the $i$th and $j$th token $t_{i}$ and $t_{j}$ is determined by the combination $f_{h}$ of position and token influence functions.
@@ C.2.2 Quantitative Semantic Analysis
Appendix D Automatic Pipeline Details
D.1 Addtional Oracle on Elastic Pattern Design
Figure 13: Examples of attention matrices from different attention heads (columns) and tasks (rows) of the Vicuna-7B model. The attention matrices were averaged over 256 data items per dataset. The same head shows a similar attention span across different tasks, explaining the robust cross-dataset generalizability of our method.
We visualize the attention matrix of the same attention heads across three additional tasks in Figure 13, as an extension of Figure 2. The consistent attention span across tasks sheds light on the strong cross-dataset generalization ability of our MoA method.
D.2 Derivation of Attention Influence
We use the first-order Taylor expansion to calculate the influence of each attention value. This approximation approach is supported by methodologies commonly employed in other LLM compression approaches (Li et al., 2024a; Shi et al., 2021; Das et al., 2023; Jiang et al., 2023).
@@ D.2 Derivation of Attention Influence
D.3 Optimization Details
D.3.1 Optimizing at Single Length
@@
D.3.2 Optimizing at Multiple Lengths
With the ability to optimize at a single length, we utilize the same framework for multi-objective MIP across various lengths. The key is to transform the multi-objective MIP problem into several single-objective MIP problems (Paria et al., 2018). We utilize the idea of epsilon-constraint method (Yv et al., 1971).